OpenStudy (anonymous):

Hey guys i need some help with a few questions please! 1)Find the degree of the monomial 6p3q2 2)Simplify (7t2 + 9) + (6t2 + 8). 3)Simplify (2z – 3)2

jimthompson5910 (jim_thompson5910):

the exponents on 6p^3q^2 are 3 and 2. Add them to get 3+2 = 5 So the degree of 6p^3q^2 is 5

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

for #2 remove the parenthesis and combine like terms

jimthompson5910 (jim_thompson5910):

one pair of like terms is 7t^2 and 6t^2 and they add to 13t^2

OpenStudy (anonymous):

For number 2 these are the possible answers a) t2 + 1 b) 13t2 + 17 c) 13t4 +17 d) 42t2 + 72

jimthompson5910 (jim_thompson5910):

do you see how 6t^2 + 7t^2 = 13t^2 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you add the like terms right?

jimthompson5910 (jim_thompson5910):

exactly

jimthompson5910 (jim_thompson5910):

the other pair of like terms are 9 and 8 they add to 9+8 = 17

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

for the last one, you would use the FOIL rule since (2z – 3)^2 = (2z – 3)(2z – 3)

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

4z2 – 9 4z2 – 9z 4z2 – 12z – 9 4z2 – 12z + 9 are the possible answers

jimthompson5910 (jim_thompson5910):

using the FOIL rule, what do you get

OpenStudy (anonymous):

nothing

jimthompson5910 (jim_thompson5910):

not quite

jimthompson5910 (jim_thompson5910):

(2z – 3)(2z – 3) = 2z(2z – 3) - 3(2z – 3) Now use the distributive property to distribute out 2z(2z – 3) - 3(2z – 3)

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

the answer is NOT choice A

jimthompson5910 (jim_thompson5910):

2z(2z – 3) - 3(2z – 3) 4z^2 - 6z - 6z + 9 ???